# GDGB: A Benchmark for Generative Dynamic Text-Attributed Graph Learning

### v1

@@ B.1 Details of TDGG and IDGG on GAG-General
The destination node selection and edge generation stages in TDGG naturally integrate traditional discriminative tasks (e.g., node retrieval, edge classification) with LLM-driven text understanding and generation, forming a DyTAG-specific generative paradigm. The pseudocode of TDGG on GAG-General is shown in Algorithm 1.
@@
The IDGG above faithfully replicates the evolutionary dynamics of real-world DyTAG through new node and edge generation over time. Although inherently challenging, this paradigm establishes a foundational direction for advancing dynamic graph generation research by simulating open-ended growth processes. The pseudocode of IDGG on GAG-General is shown in Algorithm 2. The source/destination node generation counts $R_{\text{src}}/R_{\text{dst}}$ for IDGG are shown in Table 12).
@@
1:  $\mathcal{N}_{\text{src}}^{0}=\mathcal{N}_{\text{src}}^{\text{init}},\mathcal{N}_{\text{dst}}^{0}=\mathcal{N}_{\text{dst}}^{\text{init}},\mathcal{E}^{0}=\mathcal{E}^{0}_{\text{init}}$
@@
B.2 Details of Metrics
We provide more detailed mathematical formulations and implementation specifics of our evaluation metrics in Section 4.2 as follows.
@@ B.2 Details of Metrics
  • Degree/Spectra MMD. This metric quantifies the discrepancy between the distribution of graph descriptors (e.g., degree/spectral properties) in the generated and ground-truth graphs MMD ; graphrnn ; EDGE . We compute the distribution distance, using the maximum mean discrepancy (MMD) with a radial basis function (RBF) kernel. Thus, lower MMD values indicate higher structural fidelity. Specifically, we employ a positive definite RBF kernel with a smoothing parameter $v$, defined as:
@@
    where $F(k)$ is the empirical cumulative distribution function (CDF) of the degree data, and $H(k)$ is the CDF of the best-fitting power-law model. $k_{\min}$ denotes the minimum degree value used for fitting, set to $k_{\min}=2$ to ensure robustness. The power-law exponent $\alpha$ indicates adherence to a power-law distribution, typically valid if $\alpha\in[2,3]$ powerlaw . A DyTAG is power-law valid if $D_{k}<0.15$ and the power-law exponent $\alpha\in[2,3]$.
Textual Quality Metrics. Inspired by recent role-playing agent research RoleBench ; SocialBench ; charbox , we employ the LLM-as-Evaluator framework to assess the text quality in the generated DyTAGs. Considering the targets and requirements of DyTAG generation tasks, we define five scoring criteria:
@@
Each criterion is scored on a 1–5 scale by LLM evaluators, ensuring a multidimensional assessment of textual quality. This approach outperforms embedding-based metrics like BERTScore bertscore by avoiding information loss during feature compression.
Graph Embedding Metric. To jointly evaluate structural, temporal, and textual fidelity, we extend the JL-Metric jlmetric to a better adaptation on DyTAGs with textual node feature integration to a graph embedding-based indicator. Specifically, the metric is calculated through three stages:
@@
  • Random Projection Mapping. Leveraging the Johnson-Lindenstrauss theorem for transforming varying dimensionality with bounded error guarantees jlmetric , we project high-dimensional embeddings into a consistent space with random matrices, which enables cross-graph alignment despite different node counts and interaction histories. The projected graph embedding is:
@@
Appendix C Supplementary Experimental Details
C.1 Large Language Model Baselines
@@
LLMs possess strong capabilities in language understanding and generation, demonstrating impressive performance across a wide range of natural language processing tasks. To comprehensively evaluate the performance of different models, we selected several representative LLMs for comparison in our experiments. These include open-source models: DeepSeek-R1-Distill-Qwen-32B deepseek , Llama-3-70B-Instruct llama3 , and Qwen2.5-72B-Instruct qwen2.5 . In addition, we include a closed-source model, GPT-4o-Mini gpt4omini , as a reference baseline.
C.2 Dynamic Graph Neural Network Baselines
@@
JODIE JODIE . This model is a representation learning framework for nodes in temporal networks that captures dynamic embedding trajectories based on sequences of interactions. JODIE uses two coupled recurrent neural networks to update the states of interacting entities and employs a projection operation to forecast future embedding trajectories.
@@
TGN TGN . This method models sequences of time-stamped events. By combining memory modules with graph-based operators, TGN effectively captures temporal and structural dependencies in evolving graphs. This innovative architecture enables TGNs to maintain historical context while updating node representations in real time.
@@
CAWN CAWN . This approach inductively represents temporal networks through an anonymization strategy based on sampled walks. These walks explore the causal structure of network dynamics and generate inductive node representations. A neural network encodes and aggregates the sampled walks to produce the final node embeddings.
@@
GraphMixer Graphmixer . This model is a simple yet effective architecture for temporal link prediction, composed of three key components: a multi-layer perceptron (MLP)-based link encoder to capture information from temporal links, a node encoder using neighbor mean-pooling, and an MLP-based link classifier for prediction.
@@
DyGFormer DyGLib . This method is a Transformer-based model for dynamic graph learning that leverages nodes’ historical first-hop interactions to learn meaningful representations. It introduces a neighbor co-occurrence encoding scheme to capture correlations between source and destination nodes based on their interaction histories. To handle longer sequences efficiently, DyGFormer uses a patching technique that divides each historical sequence into smaller segments, allowing the Transformer to process long-term dependencies effectively.
C.3 Dynamic Graph Generation Model Baselines
@@
DG-Gen DGGEN . This model is a generative framework for CTDG that enables assumption-free and inductive graph generation. Built as an encoder-decoder model, it learns conditional probabilities of temporal interactions to model topological evolution effectively. By relying on temporal embeddings instead of node IDs, DG-Gen supports inductive learning, allowing it to generate graphs with unseen nodes, timestamps, and edge features.
@@
VRDAG VRDAG . This model is a novel framework for simultaneously generating dynamic graph topology and node attributes in a data-driven manner. It uses a bi-flow graph encoder to preserve directional message flow and attribute information in node embeddings. A GRU-based module captures temporal dependencies, updating hidden node states across time. VRDAG also parameterizes a flexible prior distribution to sample latent variables at future timesteps, enabling it to model complex dependencies within and across topology and attribute evolution. This variational approach effectively captures both structural and attribute dynamics in evolving graphs.
C.4 Implementation Details
Implementation Details of GAG-General on TDGG and IDGG. Following GAG GAG , we maintain a memory module for each node in DyTAG to record the information of its historical interacted neighbors, effectively incorporating structural and temporal dynamics from the DyTAG. In our proposed GAG-General, the node memory module is implemented via random walks sauerwald2019random . We set the number of random walks for each node as 10, with a random walk length of 10. The node memory is capped at a maximum memory context length of 1,000. Additionally, our generative framework includes an optional memory reflection mechanism, which leverages the LLMs to distill node memories into valuable summaries, akin to the message aggregation process in GNNs kipf2016semi ; peng2024beyond . During the destination node selection process for the active source node agent, for both TDGG and IDGG, we set the number of recalled candidate destination nodes for each source node to 10. For the inference parameter configuration of benchmarked LLMs, we set the temperature as 0.8, the top-p (nucleus sampling) as 0.9, the repetition penalty as 1.1, and a maximum token limit as 2,000 to balance randomness, diversity, and redundancy suppression. With regarding to the LLM prompt design for TDGG and IDGG, we incorporates detailed node descriptions, edge descriptions, memory reflection instructions, interaction instructions, and node generation instructions in our proposed GAG-General. The example full templates on Sephroa (bipartite) and WeiboTech (non-bipartite) are provided in Appendix E. As for the LLM-as-Evaluator framework of the textual quality metric, based on the implementation of CharacterBox charbox , we use GPT as the LLM backbone for the evaluation framework. Experiments are conducted on NVIDIA A800 with 80 GB memory.
@@ C.4 Implementation Details
Implementation Details of TDGG. For the discriminative task comparisons in TDGG, following DTGB DTGB , node/edge texts are encoded using the BERT-base-uncased model BERT as initialization for each DGNN. In terms of discriminative task-specific metrics, we use Hit@1 and Hit@10 for node retrieval (link prediction) tasks. Specifically, for the node retrieval task, we set the number of negative samples to 100 for each positive sample. Notably, the Hit@1 metric can be considered a more challenging form of link prediction, as it involves a positive-to-negative sample ratio of 1:100, compared to the easier 1:1 ratio. For edge classification, we evaluate the models using weighted precision, recall, and F1 scores, following the settings in DTGB DTGB .
@@
Implementation Details of IDGG. For the comparisons with dynamic graph generation baselines in IDGG, we use the seed DyTAG with 1,000 edges as the input data for training DG-Gen and VRDAG. Additionally, we adjust the generation ratio in DG-Gen to generate 2,000 edges using its publicly available code DGGEN , whereas VRDAG does not offer explicit generation size control VRDAG . Therefore, evaluations are conducted based on the graphs actually generated by VRDAG. Similar to TDGG, following DTGB DTGB , node/edge texts are encoded using the BERT-base-uncased model BERT as initialization for each dynamic graph generation baseline.
Appendix D Supplementary Experimental Results
D.1 Results of Graph Structural Comparison
Figure 7: Left: Comparison of Degree MMD w/ and w/o text on GDGB datasets and DTGB datasets in VRDAG. Right: Comparison of Spectra MMD w/ and w/o text on GDGB datasets and DTGB datasets in VRDAG. (Same as Figure 2.)
D.2 Scalability Analysis
@@
Following GAG GAG , we adopt parallel processing techniques agentscope in implementing GAG-General to mitigate idle inference time in LLMs during DyTAG generation. Agents are grouped into distinct clusters: each active source node agent collaborates with its corresponding destination node agent, and these groups execute in parallel across CPU cores with multiple ports, enabling an efficient DyTAG generation pipeline.
@@
As shown in Figure 9, we evaluate the time costs of generating a DyTAG with 10,000 edges in TDGG and a DyTAG with 2,000 edges in IDGG, using a 1,000-edge seed graph across diverse datasets with GAG-General applying GPT as the LLM backbone. Under the parallel architecture, the TDGG task achieves an average generation time of 1.61 hours, demonstrating satisfying scalability for DyTAG generation. For IDGG, the additional computational overhead of generating new nodes increases total time, which is jointly determined by both node and edge counts. Across eight datasets, the IDGG task achieves an average generation time of 1.11 hours, indicating moderate scalability with room for further optimization. Furthermore, to rigorously assess the scalability of GAG-General, we conduct experiments on generating large-scale DyTAGs in both TDGG and IDGG tasks. As illustrated in Figure 8, for instance, generating a DyTAG with 100,000 edges in TDGG on Sephora takes 16.3 hours, while generating a DyTAG with 50,000 edges in IDGG requires 35.7 hours. Although the current time costs for large-scale DyTAG generation remain relatively high, our GDGB datasets—comprising multiple million-edge-scale DyTAGs—provide a robust foundation for future research on scalability challenges in DyTAG generation.
@@
D.3 Results of TDGG on Graph Structural Metrics
Comprehensive results on Degree MMD, Spectra MMD, $D_{k}$, $\alpha$, and power-law validity under TDGG using three additional LLM backbones (Deepseek, Llama, and Qwen) are available in Table 13, 14, and 15. The results demonstrate that the generated DyTAGs on most datasets exhibit low Degree/Spectra MMD values and satisfy the power-law validity criterion, confirming high-quality generation in TDGG with GAG-General.

### v2

@@ D.2 Details of TDGG and IDGG on GAG-General
The destination node selection and edge generation stages in TDGG naturally integrate traditional discriminative tasks (e.g., node retrieval, edge classification) with LLM-driven text understanding and generation, forming a DyTAG-specific generative paradigm. The pseudocode of TDGG on GAG-General is shown in Algorithm˜1.
@@
The IDGG above faithfully replicates the evolutionary dynamics of real-world DyTAG through new node and edge generation over time. Although inherently challenging, this paradigm establishes a foundational direction for advancing dynamic graph generation research by simulating open-ended growth processes. The pseudocode of IDGG on GAG-General is shown in Algorithm˜2. The source/destination node generation counts $R_{\text{src}}/R_{\text{dst}}$ for IDGG are shown in Table˜17).
@@
[…]
@@
D.3 Details of Metrics
We provide more detailed mathematical formulations and implementation specifics of our evaluation metrics in Section˜4.2 as follows.
@@ D.3 Details of Metrics
  • Degree/Spectra MMD. This metric quantifies the discrepancy between the distribution of graph descriptors (e.g., degree/spectral properties) in the generated and ground-truth graphs (Xiang et al., 2022; You et al., 2018; Chen et al., 2023). We compute the distribution distance, using the maximum mean discrepancy (MMD) with a radial basis function (RBF) kernel. Thus, lower MMD values indicate higher structural fidelity. Specifically, we employ a positive definite RBF kernel with a smoothing parameter $v$, defined as:
@@
    where $F(k)$ is the empirical cumulative distribution function (CDF) of the degree data, and $H(k)$ is the CDF of the best-fitting power-law model. $k_{\min}$ denotes the minimum degree value used for fitting, set to $k_{\min}=2$ to ensure robustness. The power-law exponent $\alpha$ indicates adherence to a power-law distribution, typically valid if $\alpha\in[2,3]$ (Clauset et al., 2009). A DyTAG is power-law valid if $D_{k}<0.15$ and the power-law exponent $\alpha\in[2,3]$.
Textual Quality Metrics. Inspired by recent role-playing agent research (Wang et al., 2024b; Chen et al., 2024; Wang et al., 2024a), we employ the LLM-as-Evaluator framework to assess the text quality in the generated DyTAGs. Considering the targets and requirements of DyTAG generation tasks, we define five scoring criteria:
@@
Each criterion is scored on a 1–5 scale by LLM evaluators, ensuring a multidimensional assessment of textual quality. For the LLM-as-Evaluator component, we consistently employ GPT as the underlying LLM backbone for all baselines, with hyperparameters (e.g., temperature, top‑p, and repetition penalty) kept identical to those specified in Section˜E.4. The prompts used in the LLM-as-Evaluator framework are provided in Section˜G.3. This approach outperforms embedding-based metrics like BERTScore (Zhang et al., 2020) by avoiding information loss during feature compression.
Graph Embedding Metric. To jointly evaluate structural, temporal, and textual fidelity, we extend the JL-Metric (Hosseini et al., 2025) to a better adaptation on DyTAGs with textual node feature integration to a graph embedding-based indicator. Specifically, the metric is calculated through three stages:
@@
  • Random Projection Mapping. Leveraging the Johnson-Lindenstrauss theorem for transforming varying dimensionality with bounded error guarantees (Hosseini et al., 2025), we project high-dimensional embeddings into a consistent space with random matrices, which enables cross-graph alignment despite different node counts and interaction histories. The projected graph embedding is:
@@
Appendix E Supplementary Experimental Details
E.1 Large Language Model Baselines
@@
LLMs possess strong capabilities in language understanding and generation, demonstrating impressive performance across a wide range of natural language processing tasks. To comprehensively evaluate the performance of different models, we selected several representative LLMs for comparison in our experiments. These include open-source models: DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025), Llama-3-70B-Instruct (AI@Meta, 2024), and Qwen2.5-72B-Instruct (Team, 2024). In addition, we include a closed-source model, GPT-4o-Mini (OpenAI, 2024), as a reference baseline.
E.2 Dynamic Graph Neural Network Baselines
@@
JODIE (Kumar et al., 2019). This model is a representation learning framework for nodes in temporal networks that captures dynamic embedding trajectories based on sequences of interactions. JODIE uses two coupled recurrent neural networks to update the states of interacting entities and employs a projection operation to forecast future embedding trajectories.
@@
TGN (Rossi et al., 2020). This method models sequences of time-stamped events. By combining memory modules with graph-based operators, TGN effectively captures temporal and structural dependencies in evolving graphs. This innovative architecture enables TGNs to maintain historical context while updating node representations in real time.
@@
CAWN (Wang et al., 2021a). This approach inductively represents temporal networks through an anonymization strategy based on sampled walks. These walks explore the causal structure of network dynamics and generate inductive node representations. A neural network encodes and aggregates the sampled walks to produce the final node embeddings.
@@
GraphMixer (Cong et al., 2023). This model is a simple yet effective architecture for temporal link prediction, composed of three key components: a multi-layer perceptron (MLP)-based link encoder to capture information from temporal links, a node encoder using neighbor mean-pooling, and an MLP-based link classifier for prediction.
@@
DyGFormer (Yu et al., 2023). This method is a Transformer-based model for dynamic graph learning that leverages nodes’ historical first-hop interactions to learn meaningful representations. It introduces a neighbor co-occurrence encoding scheme to capture correlations between source and destination nodes based on their interaction histories. To handle longer sequences efficiently, DyGFormer uses a patching technique that divides each historical sequence into smaller segments, allowing the Transformer to process long-term dependencies effectively.
E.3 Dynamic Graph Generation Model Baselines
@@
DG-Gen (Hosseini et al., 2024). This model is a generative framework for CTDG that enables assumption-free and inductive graph generation. Built as an encoder-decoder model, it learns conditional probabilities of temporal interactions to model topological evolution effectively. By relying on temporal embeddings instead of node IDs, DG-Gen supports inductive learning, allowing it to generate graphs with unseen nodes, timestamps, and edge features.
@@
VRDAG (Li et al., 2024). This model is a novel framework for simultaneously generating dynamic graph topology and node attributes in a data-driven manner. It uses a bi-flow graph encoder to preserve directional message flow and attribute information in node embeddings. A GRU-based module captures temporal dependencies, updating hidden node states across time. VRDAG also parameterizes a flexible prior distribution to sample latent variables at future timesteps, enabling it to model complex dependencies within and across topology and attribute evolution. This variational approach effectively captures both structural and attribute dynamics in evolving graphs.
TIGGER Gupta et al. (2022). This model is a scalable generative model for continuous-time dynamic graphs that overcomes the transductive limitations and node identity leakage of prior methods. Its inductive variant, TIGGER-I, learns a distribution over node embeddings rather than node IDs, enabling generation of graphs with unseen nodes. Notably, TIGGER-I employs a GAN-based module to jointly generate realistic features and graph structure, supporting flexible up- or down-sampling of graph size without requiring one-to-one node mapping—making it suitable for privacy-sensitive and feature-aware dynamic graph synthesis.
E.4 Implementation Details
@@
Implementation Details of TDGG. For TDGG, we set the size of the target expanded DyTAG as 10,000 edges. To ensure fairness in discriminative task comparisons, we adopt a dataset split of 1,000/1,000/8,000 (train/val/test) for DGNNs based on the implementation of DyGLib (Yu et al., 2023; Zhang et al., 2024a), and compare with the last 8,000 edges in the generated DyTAG from GAG-General. For edge classification, we evaluate the models using weighted Precision, Recall, and F1 scores (Zhang et al., 2024a). For the discriminative task comparisons in TDGG, following DTGB (Zhang et al., 2024a), node/edge texts are encoded using the BERT-base-uncased model (Devlin et al., 2019) as initialization for each DGNN. In terms of discriminative task-specific metrics, we use Hit@1 and Hit@10 for node retrieval (link prediction) tasks. Specifically, for the node retrieval task, we set the number of negative samples to 100 for each positive sample. Notably, the Hit@1 metric can be considered a more challenging form of link prediction, as it involves a positive-to-negative sample ratio of 1:100, compared to the easier 1:1 ratio.
@@
Implementation Details of IDGG. For IDGG, we set the size of the target expanded DyTAG as 2,000 edges, due to the additional stage of node generation. To ensure fair comparisons with dynamic graph generation baselines, we use the seed DyTAG with 1,000 edges as the input data for training DG-Gen and VRDAG and adjust the generation ratio in DG-Gen to generate 2,000 edges using its publicly available code (Hosseini et al., 2024), whereas VRDAG does not offer explicit generation size control (Li et al., 2024). Therefore, evaluations are conducted based on the graphs actually generated by VRDAG. As the current dynamic graph generation models’ lack of support for DyTAG’s textual features, our focus is on comparing the quality of generated graph structures and graph embeddings, rather than the texts. Similar to TDGG, following DTGB (Zhang et al., 2024a), node/edge texts are encoded using the BERT-base-uncased model (Devlin et al., 2019) as initialization for each dynamic graph generation baseline.
Implementation Details of GAG-General on TDGG and IDGG. Following GAG (Ji et al., 2025), we maintain a memory module for each node in DyTAG to record the information of its historical interacted neighbors, effectively incorporating structural and temporal dynamics from the DyTAG. In our proposed GAG-General, the node memory module is implemented via random walks (Sauerwald and Zanetti, 2019). We set the number of random walks for each node as 10, with a random walk length of 10. The node memory is capped at a maximum memory context length of 1,000. Additionally, our generative framework includes an optional memory reflection mechanism, which leverages the LLMs to distill node memories into valuable summaries, akin to the message aggregation process in GNNs (Kipf and Welling, 2016; Peng et al., 2024). During the destination node selection process for the active source node agent, for both TDGG and IDGG, we set the number of recalled candidate destination nodes for each source node to 10. For the inference parameter configuration of benchmarked LLMs, we set the temperature as 0.8, the top-p (nucleus sampling) as 0.9, the repetition penalty as 1.1, and a maximum token limit as 2,000 to balance randomness, diversity, and redundancy suppression. With regarding to the LLM prompt design for TDGG and IDGG, we incorporates detailed node descriptions, edge descriptions, memory reflection instructions, interaction instructions, and node generation instructions in our proposed GAG-General. The example full templates on Sephroa (bipartite) and WeiboTech (non-bipartite) are provided in Appendix˜G. As for the LLM-as-Evaluator framework of the textual quality metric, based on the implementation of CharacterBox (Wang et al., 2024a), we use GPT as the LLM backbone for the evaluation framework. The hyperparameters (temperature, top‑p, repetition penalty, etc.) of the LLM used in the textual quality metric are kept identical to those used in generation. Notably, although multiple LLMs serve as backbones in GAG-General for generation, all generated outputs are evaluated by the same GPT-based evaluator with fixed hyperparameters, ensuring robustness and reproducibility of the assessment. The prompts used in the LLM-as-Evaluator framework are provided in Section˜G.3. Experiments are conducted on NVIDIA A800 with 80 GB memory.
Appendix F Supplementary Experimental Results
F.1 Scalability Analysis
@@
Following GAG (Ji et al., 2025), we adopt parallel processing techniques (Gao et al., 2024) in implementing GAG-General to mitigate idle inference time in LLMs during DyTAG generation. Agents are grouped into distinct clusters: each active source node agent collaborates with its corresponding destination node agent, and these groups execute in parallel across CPU cores with multiple ports, enabling an efficient DyTAG generation pipeline.
@@
As shown in Figure˜9, we evaluate the time costs of generating a DyTAG with 10,000 edges in TDGG and a DyTAG with 2,000 edges in IDGG, using a 1,000-edge seed graph across diverse datasets with GAG-General applying GPT as the LLM backbone. Under the parallel architecture, the TDGG task achieves an average generation time of 1.61 hours, demonstrating satisfying scalability for DyTAG generation. For IDGG, the additional computational overhead of generating new nodes increases total time, which is jointly determined by both node and edge counts. Across eight datasets, the IDGG task achieves an average generation time of 1.11 hours, indicating moderate scalability with room for further optimization. Furthermore, to rigorously assess the scalability of GAG-General, we conduct experiments on generating large-scale DyTAGs in both TDGG and IDGG tasks. As illustrated in Figure˜8, for instance, generating a DyTAG with 100,000 edges in TDGG on Sephora takes 16.3 hours, while generating a DyTAG with 50,000 edges in IDGG requires 35.7 hours. Although the current time costs for large-scale DyTAG generation remain relatively high, our GDGB datasets—comprising multiple million-edge-scale DyTAGs—provide a robust foundation for future research on scalability challenges in DyTAG generation.
@@
F.2 Results of TDGG on Graph Structural Metrics
Comprehensive results on Degree MMD, Spectra MMD, $D_{k}$, $\alpha$, and power-law validity under TDGG using three additional LLM backbones (Deepseek, Llama, and Qwen) are available in Table˜18, 19, and 20. The results demonstrate that the generated DyTAGs on most datasets exhibit low Degree/Spectra MMD values and satisfy the power-law validity criterion, confirming high-quality generation in TDGG with GAG-General.